# LC-LLM: Explainable Lane-Change Intention and Trajectory Predictions with Large Language Models

### v1

@@ Abstract
To ensure safe driving in dynamic environments, autonomous vehicles should possess the capability to accurately predict the lane change intentions of surrounding vehicles in advance and forecast their future trajectories. Existing motion prediction approaches have ample room for improvement, particularly in terms of long-term prediction accuracy and interpretability. In this paper, we address these challenges by proposing LC-LLM, an explainable lane change prediction model that leverages the strong reasoning capabilities and self-explanation abilities of Large Language Models (LLMs). Essentially, we reformulate the lane change prediction task as a language modeling problem, processing heterogeneous driving scenario information in natural language as prompts for input into the LLM and employing a supervised fine-tuning technique to tailor the LLM specifically for our lane change prediction task. This allows us to utilize the LLM’s powerful common sense reasoning abilities to understand complex interactive information, thereby improving the accuracy of long-term predictions. Furthermore, we incorporate explanatory requirements into the prompts in the inference stage. Therefore, our LC-LLM model not only can predict lane change intentions and trajectories but also provides explanations for its predictions, enhancing the interpretability. Extensive experiments on the large-scale highD dataset demonstrate the superior performance and interpretability of our LC-LLM in lane change prediction task. Compared with baseline models, the performance of intention prediction can be improved by 3.1%, the root-mean-squared error (RMSE) of lateral trajectory prediction can be reduced by 19.4%, and the RMSE of longitudinal trajectory prediction can be reduced by 38.1%. To the best of our knowledge, this is the first attempt to utilize LLMs for predicting lane change behavior. Our study shows that LLMs can encode comprehensive interaction information for driving behavior understanding.
@@
Intention prediction, Trajectory prediction, Lane change, Fine-tuning, Interpretability, Large language models, Autonomous driving.
@@ I Introduction
Lane change prediction is a critical task for autonomous driving systems to ensure safe and efficient navigation on the road. Lane changing behavior, which refers to the action of a vehicle moving from one lane to another on the road, is one of the most complex driving behaviors due to intricate interactions with other vehicles. Accurately forecasting the underlying lane change intentions and future trajectories of surrounding vehicles in advance is essential for enabling autonomous vehicles to make informed decisions and take appropriate actions to avoid collisions and maintain smooth traffic flow.
@@
In the past five years, considerable advancements have been made in the field of lane change prediction. Various approaches have been explored to enhance performance in detecting lane change maneuvers [1, 2, 3]. However, at the early stage of a lane change maneuver when lateral displacement is minimal, it is essential to rely on the interaction information between the target vehicle and its surrounding vehicles for accurate predictions. Therefore, the ability to effectively model interaction information among vehicles becomes paramount. Previous methods exhibit limited capability in understanding interactive information, which results in weak long-term prediction performance. Additionally, lane change trajectory prediction has witnessed notable progress with the adoption of deep learning techniques, such as Graph Neural Networks (GNN) and Transformers, leading to competitive results [3, 4, 5]. Despite these successes, these deep learning-based approaches frequently suffer from a lack of interpretability, as they generate predictions about future behaviors without offering substantial explanations for their results. In summary, enhancing the capability to model interactions among vehicles and augmenting the interpretability of prediction results are two significant challenges in the domain of lane change prediction for autonomous driving systems.
@@
In recent years, advancements in the field of Large Language Models (LLMs) have opened new avenues for addressing challenges associated with lane change prediction. Noteworthy progress in LLMs [6, 7, 8] has showcased their robust information comprehension skills and powerful common sense reasoning abilities. Subsequent research by Mao et al. [9], Chen et al. [10] and Wu et al. [11] has successfully leveraged these LLMs to confront challenges in motion planning and perception, highlighting the considerable potential of LLMs in the domain of autonomous driving. However, current researches on lane change prediction have not yet incorporated LLMs. Given the LLMs’ profound understanding of complex driving scenarios and their extensive knowledge base, including common driving knowledge, it is rational to utilize LLMs in modeling interactions between vehicles for the enhancement of future driving behavior predictions. In addition, LLMs have demonstrated the capacity to leverage their vast knowledge base to generate explanations for their predictions [12, 13]. Inspired by these studies, there is potential to leverage LLMs for the explanation of predicted lane change intentions and future trajectories, thereby augmenting the interpretability of predictions in autonomous driving systems.
@@
In this paper, we propose a novel approach that leverages LLMs’ powerful reasoning capabilities and self-explanation capabilities to address lane change prediction challenges mentioned above. In essence, we reformulate the task of predicting lane change intentions and trajectories as a language modeling problem. To this end, we integrate heterogeneous inputs, including target vehicle state, surrounding vehicle state, map information, and Chain-of-Thought (CoT) statements, converting them into structured prompts in natural language to input into the LLM. Then, we employ a supervised fine-tuning technique to tailor the LLM specifically for the lane change prediction task. Furthermore, we integrate explanatory requirements into the prompts in the inference stage, thus enabling our fine-tuned model to generate explanations for the lane change prediction results. Benefiting from the powerful capabilities of LLM, our fine-tuned model demonstrates better performance and interpretability in lane change prediction.
@@
  • Our LC-LLM achieves explainable predictions. It not only predicts lane change intentions and trajectories but also generates explanations for the prediction results.
  • Extensive experiments on the highD dataset demonstrate that our LC-LLM outperforms all baseline models, showcasing superior performance and interpretability in lane change prediction.
@@
Figure 1: The pipeline of our LC-LLM. The green trajectory in the scenario on the left is the future trajectory of the target vehicle, which the model aims to predict. Observations are described using natural language prompts, which are input into the LLM. Supervised fine-tuning methods are then applied to fine-tune the LLM for accurate prediction of lane change intentions and trajectories. During the inference phase, prompts are designed to include explanatory requirements. As a result, the fine-tuned model is capable of predicting the target vehicle’s lane change intentions and future trajectory in the current frame, while simultaneously providing explanations for its predictions, thus improving the interpretability of the model’s outputs.
Figure 3: An example of our input prompts in the fine-tuning stage. The input prompts comprise a system message displayed in the upper text block and a user message presented in the lower text block. The user message mainly contains map information, the target vehicle’s state, the spatial relationships between the target vehicle and its surrounding vehicles, as well as constructed chain-of-thought statements.
@@ II RELATED WORK
The literature reviews provided by Huang et al. [14] and Mozaffari et al. [15] discuss various methods for predicting the future states of vehicles for autonomous driving systems. These prediction studies can be classified into the following two categories based on the output type: Intention prediction and trajectory prediction. Concurrently, LLMs have shown rapid progress in recent developments. Comprehensive overviews of LLMs applications in autonomous driving are discussed in [16].
@@ III Methodology
In this section, we introduce our LC-LLM, a model based on LLM designed for predicting lane change intentions and trajectory in autonomous driving systems. The whole pipeline of our LC-LLM is depicted clearly in Fig. 1. We reconceptualize the task of predicting intentions and trajectory as a language modeling problem. To this end, we articulate observations using natural language as prompts for input into the LLM and leverage supervised fine-tuning techniques to tailor the LLM to this specific task. During the inference stage, we incorporate explanatory requirements into the prompt. As a result, our fine-tuned model, LC-LLM, not only forecasts the lane change intentions and future trajectories but also provides explanations for the predictions, thereby enhancing their interpretability.
@@ III-A Problem Formulation
We define the state of the target vehicle as tv, the states of the surrounding vehicles as sv, and the map information as m, which serve as the inputs to our model. The outputs are the lane change intention $\mathcal{I}$ and the future trajectory $\mathcal{T}$. The process of the model $\mathcal{F}$ can be formulated as follows:
$\mathcal{I},\mathcal{T}=\mathcal{F}(\textbf{{tv}},\textbf{{sv}},\textbf{{m}})$ (1)
The $\mathcal{I}\in\{0,1,2\}$, where 0 denotes “keep lane”, 1 denotes ”left lane change”, and 2 denotes ”right lane change”. The future trajectory of prediction horizon $t$ can be denoted by $\mathcal{T}=\{(\textit{x}_{1},\textit{y}_{1}),(\textit{x}_{2},\textit{y}_{2}),...,(\textit{x}_{t},\textit{y}_{t})\}$. The target vehicle state tv consists of a historical trajectory of this vehicle and its current velocity and type. The surrounding vehicle states m contain each vehicle’s current velocity, type, and their spatial relationship relative to the target vehicle. The map information m includes details such as lane identifiers and lane markings.
In this paper, we reframe the task of predicting intentions and trajectory as a language modeling problem. By utilizing natural language to describe both the input data and prediction results, we can represent them as a sequence of tokens. The input data which describe the current frame driving scenario can be denoted as a sequence of tokens $\textit{T}_{s}$:
@@
$\{\textit{T}_{1},\textit{T}_{2},...,\textit{T}_{n}\}=K(\{\mathcal{I},(\textit{x}_{1},\textit{y}_{1}),(\textit{x}_{2},\textit{y}_{2}),...,(\textit{x}_{t},\textit{y}_{t})\})$ (3)
@@ III-B Prompting
As illustrated in Fig. 3, the input prompts consist of a system message displayed in the upper text block and a user message presented in the lower text block. The system message maintains consistency across diverse driving scenarios. It delineates the designated role of the LLM, provides details of the coordinate system, and outlines the information and format for the LLM’s output. In our paper, the assigned role for the LLM is that of a predictive model integrated within an autonomous driving system. The coordinate system in each driving scenario is the vehicle coordinate system, which is centered on the current position of the target vehicle. The expected output includes predictions of lane change intentions and trajectory points over a future time horizon of four seconds.
@@
The user message provides a description of the observations specific to the current frame, thus varying with each driving scenario. It includes information about the map, the state of the target vehicle, the spatial relationships between the target vehicle and its surrounding vehicles, as well as constructed Chain-of-Thought (CoT) [34] statements. Map information in the prompts primarily denotes the number of lanes in the scenario and indicates whether the target vehicle is located in the leftmost, middle, or rightmost lane. Prompts related to the target vehicle’s state are generated by detailing the target vehicle’s historical trajectory over the past two seconds, its current velocity, and its vehicle type. Given that most vehicles do not have a large lateral displacement 4 seconds before changing lanes, the model mainly relies on the interaction information between the target vehicle and surrounding vehicles to predict lane change intentions in advance. Therefore, comprehending the information pertaining to surrounding vehicles is crucial for accurately predicting the lane change intentions of the target vehicle. We denote the information of the nearest vehicles in eight directions surrounding the target vehicle as surrounding vehicle prompts. These directions include the front, left front, right front, left side, right side, rear, left rear, and right rear. The surrounding vehicle information in each direction encompasses details such as vehicle type, current speed, and relative distance from the target vehicle. This comprehensive surrounding vehicle information serves as a critical prompt for the predictive model, enabling a thorough analysis of the contextual interactions influencing the target vehicle’s lane change intentions.
Recently, CoT prompting has demonstrated remarkable capabilities in carrying out more intricate reasoning tasks [35, 34, 36]. Drawing inspiration from Zero-shot-CoT [34] and Least-to-Most prompting, we formulate CoT prompts by subdividing the task into three subproblems. These subproblems serve as guidance for the LLM in the reasoning process, facilitating the subsequent resolution of the final prediction problem. In the aforementioned prompt, a wealth of information is provided to the LLM. However, the LLM may not thoroughly comprehend this information or may resort to simplistic problem-solving strategies. Therefore, the incorporation of CoT prompts is deemed necessary to instruct the LLM to deliberate thoughtfully on aspects such as target vehicle information, potential intentions of surrounding vehicles, and the spatial relationships and interactions between the target vehicle and its surroundings. The details of CoT prompts are shown in Fig. 3.
The last portion of our prompts details the content and format of the final answer. Our model’s objective is to accurately forecast lane change intentions and trajectories. Consequently, the final answer should encapsulate two tasks: intention prediction and trajectory prediction. It is noted in the literature that high-level intention prediction substantially contributes to trajectory prediction [22, 3]. Therefore, aligning with the characteristics of the autoregressive language model, we have formulated the final answer to sequentially output the intention first, followed by the corresponding future trajectory points.
@@
Figure 5: An example of our input prompts in the inference stage. The content in the green text block is the added prompts in the inference stage. The ellipsis part refers to other content of the input prompts, the same as Fig. 3.
III-C Fine-tuning

### v2

@@ Abstract
To ensure safe driving in dynamic environments, autonomous vehicles should possess the capability to accurately predict lane change intentions of surrounding vehicles in advance and forecast their future trajectories. Existing motion prediction approaches have ample room for improvement, particularly in terms of long-term prediction accuracy and interpretability. In this paper, we address these challenges by proposing LC-LLM, an explainable lane change prediction model that leverages the strong reasoning capabilities and self-explanation abilities of Large Language Models (LLMs). Essentially, we reformulate the lane change prediction task as a language modeling problem, processing heterogeneous driving scenario information as natural language prompts for LLMs and employing supervised fine-tuning to tailor LLMs specifically for lane change prediction task. Additionally, we finetune the Chain-of-Thought (CoT) reasoning to improve prediction transparency and reliability, and include explanatory requirements in the prompts during inference stage. Therefore, our LC-LLM model not only predicts lane change intentions and trajectories but also provides CoT reasoning and explanations for its predictions, enhancing its interpretability. Extensive experiments based on the large-scale highD dataset demonstrate the superior performance and interpretability of our LC-LLM in lane change prediction task. To the best of our knowledge, this is the first attempt to utilize LLMs for predicting lane change behavior. Our study shows that LLMs can effectively encode comprehensive interaction information for driving behavior understanding.
@@
Lane change, Large language models, Intention prediction, Trajectory prediction, Fine-tuning, Interpretability, Autonomous driving.
@@ I Introduction
Lane change prediction is a critical task for autonomous driving systems to ensure safe and efficient navigation on the road. Lane changing behavior, which refers to the action of a vehicle moving from one lane to another, is one of the most complex driving behaviors due to intricate interactions with other vehicles. Accurately forecasting the underlying lane change intentions and future trajectories of surrounding vehicles in advance is essential for autonomous vehicles to make informed decisions and take appropriate actions.
@@
In recent years, considerable advancements have been made in lane change predictions. Various approaches have been explored to enhance performance in detecting lane change maneuvers [1, 2, 3, 4]. However, at the early stage of a lane change maneuver when lateral displacement is minimal, it is essential to rely on the interaction information between the target vehicle and its surrounding vehicles for accurate predictions. Therefore, the ability to effectively model interaction information among vehicles becomes paramount. Previous methods exhibit limited capability in understanding interactive information, which results in weak long-term prediction performance [1, 3]. Additionally, lane change trajectory prediction has witnessed notable progress with the adoption of deep learning techniques, such as Graph Neural Networks (GNN) and Transformers, leading to competitive results [4, 5, 6]. Despite these successes, these deep learning-based approaches frequently suffer from a lack of interpretability, as they generate predictions about future behaviors without offering substantial explanations for their results. In summary, enhancing the capability to model interactions among vehicles and augmenting the interpretability of prediction results are key challenges in lane change prediction for autonomous driving.
@@
In recent years, advancements in Large Language Models (LLMs) have provided new opportunities for addressing challenges associated with lane change prediction. Noteworthy progress in LLMs [7, 8, 9] has showcased their robust information comprehension skills and powerful common sense reasoning abilities. However, current researches on lane change prediction have not yet incorporated LLMs. Given the LLMs’ profound understanding of complex driving scenarios and their extensive knowledge base, including common driving knowledge, it is rational to utilize LLMs in modeling interactions between vehicles for the enhancement of future driving behavior predictions. In addition, LLMs have demonstrated the capacity to leverage their vast knowledge base to generate explanations for their predictions [10, 11]. Inspired by these studies, there is potential to leverage LLMs for the explanation of predicted lane change intentions and future trajectories, thereby augmenting the interpretability of predictions in autonomous driving.
@@
In this paper, we propose a novel approach that leverages LLMs’ powerful reasoning and self-explanation capabilities to address lane change prediction challenges mentioned above. In essence, we reformulate the task of predicting lane change intentions and trajectories as a language modeling problem. To this end, we integrate heterogeneous inputs, including target vehicle state, surrounding vehicle state, and map information, converting them into structured prompts in natural language to input into the LLM. Then, we employ a supervised fine-tuning technique to tailor the LLM specifically for the lane change prediction task. Additionally, in fine-tuning stage, we perform Chain-of-Thought (CoT) reasoning to enhance transparency and reliability in lane change predictions. Finally, we integrate explanatory requirements into the prompts in the inference stage, thus enabling our fine-tuned model to generate explanations for the lane change prediction results. Benefiting from the powerful capabilities of LLM, our fine-tuned model demonstrates better performance and interpretability in lane change prediction.
@@
  • Our LC-LLM achieves explainable predictions. It not only predicts lane change intentions and trajectories but also generates explanations for the prediction results. Additionally, we evaluate the accuracy of CoT reasoning to quantitatively evaluate the interpretability of our LC-LLM.
  • Extensive experiments on the highD dataset demonstrate that our LC-LLM outperforms all baseline models, achieving a 17.7% improvement in lane change intention prediction, a 64.4% improvement in lateral trajectory prediction, and a 66.1% improvement in longitudinal trajectory prediction, respectively.
@@ II RELATED WORK
The literature reviews [12, 13, 14] discuss various methods for predicting the future states of vehicles for autonomous driving systems. These prediction studies can be classified into the following two categories based on the output type: Intention prediction and trajectory prediction. Concurrently, LLMs have shown rapid progress in recent developments. Comprehensive overviews of LLMs applications in autonomous driving are discussed in [15].
@@ II-A Intention Prediction
Figure 1: The pipeline of our LC-LLM. The green trajectory in the Lane Change Event is the future trajectory of the target vehicle, which the model aims to predict. Observations are described using natural language prompts, which are input into the LLM. Supervised fine-tuning methods are then applied to fine-tune the LLM for accurate prediction of lane change intentions and trajectories, as well as the CoT reasoning. During the inference phase, prompts are designed to include explanatory requirements. As a result, the fine-tuned model is capable of predicting the target vehicle’s lane change intentions and future trajectory in the current frame, while simultaneously providing the thought reasoning and explanations for its predictions, thus improving the interpretability of the model’s outputs.
@@ III Methodology
In this section, we introduce our LC-LLM, a model based on LLM designed for predicting lane change intentions and trajectory in autonomous driving systems. The whole pipeline of our LC-LLM is depicted clearly in Fig. 1. We reconceptualize the task of predicting intentions and trajectory as a language modeling problem. To this end, we articulate observations using natural language as prompts for input into the LLM and leverage supervised fine-tuning techniques to tailor the LLM to this specific task. During the inference stage, we incorporate explanatory requirements into the prompt. As a result, our fine-tuned model, LC-LLM, not only forecasts the lane change intentions and future trajectories but also provides CoT reasoning and explanations for the predictions, thereby enhancing their interpretability.
Figure 3: An example of our input prompts in the fine-tuning stage. The input prompts comprise a system message displayed in the upper text block and a user message presented in the lower text block. The user message mainly contains map information, the target vehicle’s state, the spatial relationships between the target vehicle and its surrounding vehicles.
@@ III-A Problem Formulation
We define the state of the target vehicle as tv, the states of the surrounding vehicles as sv, and the map information as m, which serve as the inputs to our model. The outputs are the CoT reasoning $\mathcal{C}$, the lane change intention $\mathcal{I}$ and the future trajectory $\mathcal{T}$. The process of the model $\mathcal{F}$ can be formulated as follows:
$\mathcal{C},\mathcal{I},\mathcal{T}=\mathcal{F}(\textbf{{tv}},\textbf{{sv}},\textbf{{m}})$ (1)
The $\mathcal{I}\in\{0,1,2\}$, where 0 denotes “keep lane”, 1 denotes “left lane change”, and 2 denotes “right lane change”. The future trajectory of prediction horizon $t$ can be denoted by $\mathcal{T}=\{(\textit{x}_{1},\textit{y}_{1}),(\textit{x}_{2},\textit{y}_{2}),...,(\textit{x}_{t},\textit{y}_{t})\}$. And the corresponding CoT reasoning $\mathcal{C}$ is composed of notable features and potential behavious. The target vehicle state tv consists of a historical trajectory of this vehicle and its current velocity, type and so on. The surrounding vehicle states m contain each vehicle’s current velocity, type, and their spatial relationship relative to the target vehicle. The map information m includes details such as lane identifiers and lane markings.
We reframe the task of predicting intentions and trajectory as a language modeling problem. By utilizing natural language to describe both the input data and output results, we can represent them as a sequence of tokens. The input data which describe the current frame driving scenario can be denoted as a sequence of tokens $\textit{T}_{s}$:
@@
$\{\textit{T}_{1},\textit{T}_{2},...,\textit{T}_{n}\}=K(\mathcal{C},\mathcal{I},\mathcal{T})$ (3)
@@ III-B Prompting
As illustrated in Fig. 3, the input prompts consist of a system message displayed in the upper text block and a user message presented in the lower text block. The system message maintains consistency across diverse driving scenarios. It delineates the designated role of the LLM, provides details of the coordinate system, and outlines the information and format for the LLM’s output. In our paper, the assigned role for the LLM is that of a predictive model integrated within an autonomous driving system. The coordinate system in each driving scenario is the vehicle coordinate system, which is centered on the current position of the target vehicle. The expected output includes predictions of lane change intentions and trajectory points over a future time horizon of four seconds, as well as includes thought reasoning which consists of notable features and potential behavior.
@@
The user message provides a description of the observations specific to the current frame, thus varying with each driving scenario. It includes information about the map, the state of the target vehicle, the spatial relationships between the target vehicle and its surrounding vehicles. Map information in the prompts primarily denotes the number of lanes in the scenario and indicates whether the target vehicle is located in the leftmost, middle, or rightmost lane. Prompts related to the target vehicle’s state are generated by detailing the target vehicle’s historical trajectory over the past two seconds, its current velocity, and its vehicle type. Given that most vehicles do not have a large lateral displacement 4 seconds before changing lanes, the model mainly relies on the interaction information between the target vehicle and surrounding vehicles to predict lane change intentions in advance. Therefore, comprehending the information pertaining to surrounding vehicles is crucial for accurately predicting the lane change intentions of the target vehicle. We denote the information of the nearest vehicles in eight directions surrounding the target vehicle as surrounding vehicle prompts. These directions include the ahead, left front, right front, left side, right side, rear, left rear, and right rear. The surrounding vehicle information in each direction encompasses details such as vehicle type, current speed, and relative distance from the target vehicle. This comprehensive surrounding vehicle information serves as a critical prompt for the predictive model, enabling a thorough analysis of the contextual interactions influencing the target vehicle’s lane change intentions.
Figure 4: An example of the output of our LC-LLM. The content in the yellow text block is the Chain-of-Thought reasoning.
III-C Reasoning
Recently, CoT reasoning has demonstrated remarkable capabilities in carrying out more intricate reasoning tasks [30, 35, 36, 37]. Drawing inspiration from GPT-Driver [30], which utilizes Chain-of-Thought reasoning strategy in motion planning tasks to enhance transparency throughout the planning procedure, we employ CoT reasoning in lane change prediction task to improve reliability of predictions and explanations. Unlike GPT-Driver, which only labels thoughts based on whether the ego vehicle and surrounding vehicles will collide in the future, we also consider domain knowledge in driving, traffic rules, and traditional lane change model rules [38, 39] for labeling CoT reasoning. In this way, our LC-LLM model can learn this knowledge and rules to provide more reliable and accurate predictions and explanations. The details of CoT reasoning are shown in Fig. 4.
Our CoT reasoning consists of notable features and potential behaviors. The annotation of CoT reasonings are as follows:
III-C1 Labeling of Notable feature
Firstly, we labeled notable features include significant lateral movement when lateral velocity exceeds $1.5km/h$, and high longitudinal acceleration when it surpasses $0.4m/s^{2}$. Secondly, the relative speeds of surrounding vehicles compared to the target vehicle in the directions ahead, left front, and right front are considered notable features. Vehicles in these positions with a speed higher than the target vehicle are labeled as free. Conversely, vehicles with a speed lower than the target vehicle are labeled as blocked, indicating potential congestion or the need for the target vehicle to overtake. At last, the presence of a truck ahead within 100 meters is also notable due to its impact on traffic flow and overtaking maneuvers [38]. Additionally, in right lane change scenarios, the characteristic of the target vehicle being a truck is also labeled as notable due to the traffic rule that trucks typically travel in the right lane.
III-C2 Labeling of Potential behavior
Potential behaviors are classified into eight categories: “Change to the left lane for overtaking” occurs when encountering a slower vehicle ahead and occupying the rightmost or middle lane; “Change left to the fast lane” corresponds to significant acceleration from the target vehicle; “Irregular left lane change” applies to scenarios in the highD dataset that involve a left lane change but do not fit the first two categories; “Change to the right lane for overtaking” is when the target vehicle occupies the leftmost or middle lane and the lane ahead is blocked; “Change right to the slow lane” is associated with significant deceleration from the target vehicle or when the target vehicle is a truck; “Irregular right lane change” covers right lane change scenarios in the highD dataset that do not fit the first two rightward categories; “Following and keep lane” occurs when following a slower vehicle ahead; “Normal keep lane” is when the lane ahead is unobstructed and it is a lane keeping scenario in the highD dataset. These categories assist in predicting driving patterns and providing reliable explanations.
@@ III-C2 Labeling of Potential behavior
III-D Fine-tuning